# Towards Inclusive Communication: A Unified LLM-Based Framework for Sign Language, Lip Movements, and Audio Understanding

### v1

@@ 3.3 Modality Alignment & Fusion
Concatenation & Mapping to Unified Linguistic Tokens
@@
3.4 Multi‑Task Training Strategy
Task‑Adaptive Modality Masking
@@ 3.4 Multi‑Task Training Strategy
Two‑Stage Multi-Task Training
@@
4 Experimental Setup
4.1 Dataset
@@
How2Sign (Duarte et al. 2021) is a widely used American Sign Language (ASL) dataset for the SLT task. It comprises 80 hours of instructional videos covering 10 different topics. The manually re-aligned video clips are used for training and evaluation, consisting of 31,128 sentences for training, 1,741 for validation, and 2,322 for testing.
@@
LRS3 (Afouras, Chung, and Zisserman 2018) is a widely used English dataset for ASR, VSR, and AVSR. It comprises approximately 433 hours of TED Talk videos with aligned audio and human-annotated transcriptions. The dataset contains 151K utterances for training and 1,321 utterances for testing.
4.2 Evaluation Metrics
@@
To evaluate VSR, ASR, and AVSR performance, we employ Word Error Rate (WER), a widely used metric in prior research (Afouras et al. 2018; Ma et al. 2023). A lower WER reflects better recognition accuracy. For SLT evaluation, inspired by prior studies (Rust et al. 2024), we utilize BLEU-4 (Papineni et al. 2002), ROUGE (Lin 2004), and BLEURT (Sellam, Das, and Parikh 2020), where higher scores correspond to better translation quality.
4.3 Implementation details
Videos and audio are preprocessed at 25 FPS and 16 kHz, respectively. Signer regions are cropped following (Jang et al. 2025b), and lip regions are extracted using a RetinaFace (Deng et al. 2020) facial landmark detector. We employ 1D convolutions as length adapters to temporally downsample all feature sequences to a common resolution of 12.5 Hz. We use LLaMA 3.2-3B (Touvron et al. 2023) as the LLM decoder. We employ a two-layer linear mapping network to transform the concatenated features into 3072-dimensional tokens before feeding them into the LLM. Further details on preprocessing, model dimensions, and training hyperparameters are provided in the Appendix.
Training and Evaluation
@@ 4.3 Implementation details
During training, we apply data augmentation for the VSR, AVSR tasks by randomly cropping mouth patches to $88{\times}88$ and applying horizontal flips. For SLT, we use text-level augmentation by randomly dropping 0–20% of the words from the ground-truth target sentences, following (Jang et al. 2025b). To train the first stage on VSR and SLT, we use a tri-stage scheduler with 18K warmup steps, 36K decay steps, and a peak learning rate of 1e-4. The second stage jointly trains ASR, VSR, AVSR, and SLT using the same scheduler structure, but with 500 warmup steps and 29.5K decay steps. During this stage, we apply babble noise from the NOISEX dataset (Varga and Steeneken 1993) to audio inputs with 75% probability when training ASR and AVSR tasks. The signal-to-noise ratio (SNR) is uniformly sampled from {-5, 0, 5, 10, 15, 20} dB. Training is conducted on 8 NVIDIA A6000 GPUs (48GB each), with a maximum of 1,250 frames processed per GPU per batch. For evaluation, we use beam search (width 5, temperature 0.3) for ASR, VSR, and AVSR, and greedy decoding for SLT.
@@
SSLT (2024) - - 55.7
Youtube-ASL (2023) 12.4 - -
Uni-Sign (2025) 14.9 36.0 49.4
SSVP-SLT (2024) 15.5 38.4 49.6
SSLT (2024) - - 34.0
SSVP-SLT (2024) 7.0 25.7 39.3
Fla-LLM (2024) 9.7 27.8 -
VAP (2024) 12.9 27.8 -
LiTFiC (2025b) 12.7 32.5 45.3
PGG-SLT (2025) 13.7 32.9 -
@@
Table 1: Performance comparison of SLT models on the How2Sign. $\dagger$ indicates models that are pretrained on a larger ASL corpus, YouTube-ASL (Uthus, Tanzer, and Georg 2023), which contains 984 hours of video. “AV Input” indicates whether the model handles audio or lip movement modalities.
@@
V2P (2018) 3,886 55.1
RNN-T (2019) 31,000 33.6 4.8 4.5
VTP (2022) 2,676 30.7
Auto-AVSR (2023) 1,902 23.5 1.0 1.0
Auto-AVSR (2023) 3,448 19.1 1.0 0.9
ViT3D-CM (2022) 90,000 17.0 1.6
SynthVSR (2023) 6,720 16.9
LP Conf (2024) 100,000 12.8 0.9
@@
AV-HuBERT (2022) 1,759 26.9
RAVEn (2022) 1,759 23.1 1.4
USR (2024a) 1,759 22.3 1.2 1.1
AV-HuBERT (2022) 433 28.6 1.3 1.4
VATLM (2023) 433 28.4 1.2
RAVEn (2022) 433 28.2 1.4
BRAVEn (2024b) 433 26.6 1.2
u-HuBERT (2022) 433 29.1 1.5 1.3
@@
Llama-AVSR (2025) 1,759 24.0 0.79 0.77
MMS-Llama (2025) 1,759 0.72
VSP-LLM (2024a) 433 25.4
Llama-AVSR (2025) 433 25.3 1.1 0.95
MMS-Llama (2025) 433 0.90
@@
Table 2: Performance comparison of VSR, ASR, and AVSR models on LRS3. $\dagger$ indicates models that leverage pseudo-labels generated from the VoxCeleb2 (2018) dataset (1,326 hours) during fine-tuning. SL denotes sign language.
5 Experimental Results
5.1 Comparison with the state-of-the-art methods
In Table 1, we compare the SLT performance of our unified framework against prior methods on the How2Sign dataset. Compared to PGG-SLT (Guo, Li, and Cohn 2025), the previous best model without using external data, our approach improves BLEU-4 by +1.5 (15.2 vs. 13.7) and ROUGE by +3.3 (36.2 vs. 32.9). It also outperforms LiTFiC (Jang et al. 2025b) by +1.8 BLEURT (47.1 vs. 45.3). These results highlight the advantage of jointly modeling SLT and VSR within a single framework.
In Table 2, we compare performance on VSR, ASR, and AVSR using the LRS3 dataset. Since our model is LLM-based, we focus on comparisons with recent LLM-based approaches, all of which are task-specific. For VSR, our method achieves a WER of 25.9%, comparable to task-specific models such as VSP-LLM and Llama-AVSR. In ASR, our model reaches a WER of 0.79% using only 433 hours of training data, matching the performance of Llama-AVSR, which was trained on 1,759 hours. For AVSR, our model obtains a WER of 0.93%, close to MMS-Llama (0.90%). Compared to USR (Haliassos et al. 2024a), a recent unified model for VSR, ASR, and AVSR, our method performs better on ASR (0.79% vs. 1.2%) and AVSR (0.93% vs. 1.1%) while using significantly less training data. These results demonstrate that a single unified model can effectively process sign language, lip movements, and audio signals, achieving competitive or superior performance compared to task-specific models across all four tasks. Moreover, the strong performance across tasks suggests that shared linguistic representations and cross-modal learning can generalize effectively across modalities.
@@ 5.1 Comparison with the state-of-the-art methods
Table 3: Ablation study on SLT performance on How2Sign, with progressively added lip encoder and VSR loss.
Figure 2: Training progress of the multi-task model across AVSR, ASR, VSR, and SLT tasks. Each curve shows the next-token prediction accuracy over epochs.
@@
Table 4: Effect of our two-stage multi-task training strategy. Stage 1 trains only on visual tasks (VSR, SLT), while Stage 2 jointly trains all tasks.
@@
Table 5: Effect of modality-specific dropout probabilities on multi-task learning performance.
5.2 Ablation study
Effectiveness of Lip-Aware SLT via VSR Modeling
To assess the effectiveness of our lip-aware SLT framework, which incorporates auxiliary VSR modeling within a unified architecture, we conduct controlled ablation experiments by incrementally adding each component. Four configurations are evaluated, as summarized in Table 3.
The base configuration includes the sign encoder, sign length adapter, and LLM decoder, trained solely on the SLT task using the How2Sign dataset. This setup yields a BLEU-4 score of 12.4, ROUGE of 32.6, and BLEURT of 44.6. We then add a lip encoder that processes the signer’s mouth region and injects this representation into the LLM alongside sign features. This addition leads to substantial gains: +1.5 BLEU-4, +2.4 ROUGE, and +1.6 BLEURT (13.9 / 35.0 / 46.2), highlighting the importance of explicitly modeling non-manual cues such as lip movements. Finally, we incorporate a VSR loss to train our framework by leveraging an external LRS3 dataset featuring non-signers. This additional supervision further improves the scores to 14.4 BLEU-4, 35.6 ROUGE, and 46.4 BLEURT. We attribute this improvement to the auxiliary task encourages the LLM to internalize the grammar, vocabulary, and contextual semantics of the spoken target language, thereby enhancing its ability to generate well-formed translations from visual sign inputs. This aligns with findings from recent work (Zhang, Tanzer, and Firat 2024), which showed that jointly training machine translation with SLT improves overall translation performance.
Effectiveness of Two-stage Multi-Task Learning
As shown in Figure 2, preliminary multi-task training experiments revealed significant convergence imbalances: audio-based tasks (ASR, AVSR) quickly surpassed 90% next-token prediction accuracy within 2–3 epochs, while visual tasks (VSR, SLT) improved more slowly. Motivated by these observations, we propose a two-stage multi-task learning strategy. To evaluate the effectiveness of this approach, we conducted three configurations: (1) all tasks trained jointly from scratch; (2) only Stage 1 training; and (3) full two-stage training. Results are summarized in Table 4.
@@ 5.2 Ablation study
Impact of Modality Dropout on Multi-Task Learning
Since the LRS3 dataset provides paired audio-visual inputs, dropping the lip modality corresponds to performing ASR, while dropping the audio modality corresponds to performing VSR. Based on this fact, we leverage modality dropout (Shi et al. 2022) as an implicit mechanism to control the sampling ratio of each task during training.
@@
6 Conclusion
Appendix A Training Details
Pre-processing
All videos are resampled to 25 FPS, and audio signals are resampled to 16 kHz. For sign language videos, signer regions are cropped following (Jang et al. 2025b). Mouth regions are extracted using a RetinaFace-based facial landmark detector (Deng et al. 2020), producing $96{\times}96$ patches. For the LRS3 dataset, we apply the same preprocessing to lip videos and extract log-Mel spectrograms from the audio. The resulting video crops and audio spectrograms are used as inputs to their respective encoders.
Architecture Overview
The sign encoder is based on the Video Swin-T model (Liu et al. 2022), fine-tuned for isolated sign language recognition on the How2Sign dataset by (Jang et al. 2025b). It produces 25 feature vectors (one per frame) with a dimensionality of 768. The lip encoder utilizes AV-HuBERT Large (Shi et al. 2022), which is pre-trained with pseudo-labels derived from audio-visual speech units on the LRS3 and VoxCeleb2 datasets (Chung, Nagrani, and Zisserman 2018). It outputs 25 feature vectors per second, each with a dimensionality of 1024. The audio encoder is Whisper-Medium (Radford et al. 2023), which is pre-trained and fine-tuned on large-scale supervised audio data. It encodes 50 feature vectors per second, each with a dimension of 1024. To unify the temporal resolution across modalities, we employ modality-specific 1D convolutional length adapters to downsample all feature sequences to 12.5 Hz. Specifically, for the audio features (originally at 50 Hz), we use a 1D convolution with a kernel size and stride of 4. For the lip and sign features (25 Hz), we use a kernel size and stride of 2. The downsampled features are concatenated (2816-dimensional vector) and passed through a two-layer linear projection network, with an intermediate dimension of 2944 (the average of input and output) and an output dimension of 3072 to match the LLaMA 3.2-3B (Touvron et al. 2023) embedding size. We employ LLaMA 3.2-3B as the decoder and fine-tune it using LoRA (Hu et al. 2021), applied to the query, key, value, and output projection layers. LoRA uses a rank of 16 and a scaling factor of 32.
Training Configuration
We apply data augmentation to the VSR and AVSR tasks by randomly cropping mouth patches to $88{\times}88$ and applying horizontal flipping. For the SLT task, we adopt text-level augmentation by randomly dropping 0–20% of the words in the ground-truth target sentences, following (Jang et al. 2025b).
Training proceeds in two stages: Stage 1: Train the model on VSR and SLT tasks using a tri-stage learning rate scheduler with 18K warm-up steps, 36K decay steps, and a peak learning rate of 1e-4. Stage 2: Jointly train ASR, VSR, AVSR, and SLT tasks with the same scheduler structure but 500 warm-up steps and 29.5K decay steps. During Stage 2, we apply babble noise from the NOISEX dataset (Varga and Steeneken 1993) to the audio input with a probability of 75% for the ASR and AVSR tasks. The signal-to-noise ratio (SNR) is uniformly sampled from the set $\{-5,0,5,10,15,20\}$ dB. Training is conducted on 8 NVIDIA A6000 GPUs (48GB each), with a maximum of 1,250 video frames processed per GPU per batch. A fixed random seed of 1 is used across all experiments to ensure reproducibility. We save two checkpoints during training: (i) the best checkpoint based on next-token prediction accuracy on the validation set, and (ii) the last checkpoint at the end of training. For each experiment, we compare both checkpoints and select the final model based on overall performance on BLEU-4 (for SLT) and WER (for ASR, VSR, and AVSR). All experiments are conducted with a single run due to computational constraints. For evaluation, we use beam search (width 5, temperature 0.3) for ASR, VSR, and AVSR tasks, and greedy decoding for SLT.
Appendix B Additional Results
@@
Table 6: Performance comparison between task-specific training and joint multi-task training using our proposed framework. Each of the first four rows shows the result when the model is trained on a single task only.

### v2

@@ III-C1 Length Adapters
III-C2 Concatenation & Mapping to Unified Linguistic Tokens
@@
III-D Multi‑Task Training Strategy
III-D1 Task‑Adaptive Modality Masking
@@
III-D2 Two‑Stage Multi-Task Training
@@
IV Experimental Setup
IV-A Dataset
@@
How2Sign [58] is a widely used American Sign Language (ASL) dataset for the SLT task. It comprises 80 hours of instructional videos covering 10 different topics. The manually re-aligned video clips are used for training and evaluation, consisting of 31,128 sentences for training, 1,741 for validation, and 2,322 for testing.
@@
LRS3 [35] is a widely used English dataset for ASR, VSR, and AVSR. It comprises approximately 433 hours of TED Talk videos with aligned audio and human-annotated transcriptions. The dataset contains 151K utterances for training and 1,321 utterances for testing.
IV-B Evaluation Metrics
@@
To evaluate VSR, ASR, and AVSR performance, we employ Word Error Rate (WER), a widely used metric in prior research [15, 45]. A lower WER reflects better recognition accuracy. For SLT evaluation, inspired by prior studies [6], we utilize BLEU-4 [66], ROUGE [67], and BLEURT [68], where higher scores correspond to better translation quality.
IV-C Implementation details
Videos are resampled to 25 FPS and audio signals to 16 kHz. For sign language videos, signer regions are cropped following [65], and lip regions are extracted using a RetinaFace-based facial landmark detector [69], producing $96{\times}96$ patches. For the LRS3 dataset, we apply the same preprocessing to lip videos and extract log-Mel spectrograms from the audio.The sign encoder is based on Video Swin-T [55], fine-tuned on the How2Sign dataset [65], producing 25 frame-level features of dimension 768. The lip encoder uses AV-HuBERT Large [43], pre-trained on LRS3 and VoxCeleb2 [70], outputting 25 feature vectors per second of dimension 1024. The audio encoder is Whisper-Medium [31], producing 50 features per second of dimension 1024. To unify temporal resolution, we employ modality-specific 1D convolutions to downsample all features to 12.5 Hz: kernel size/stride 4 for audio and 2 for lip and sign features. The resulting features (2816-dim) are concatenated and passed through a two-layer linear projection (intermediate dimension 2944, output 3072) to match the embedding size of LLaMA 3.2-3B [34]. We use LLaMA 3.2-3B as the decoder, fine-tuned with LoRA [71] (rank 16, scaling factor 32) applied to the query, key, value, and output projection layers.
IV-D Training and Evaluation
@@
We apply data augmentation to the VSR and AVSR tasks by randomly cropping mouth patches to $88{\times}88$ and applying horizontal flipping. For the SLT task, we adopt text-level augmentation by randomly dropping 0–20% of the words in the ground-truth target sentences, following [65]. During Stage 2 training, we further apply babble noise from the NOISEX dataset [72] to the audio input with a probability of 75% for the ASR and AVSR tasks, where the signal-to-noise ratio (SNR) is uniformly sampled from $\{-5,0,5,10,15,20\}$ dB. Training proceeds in two stages: Stage 1 trains the model on VSR and SLT tasks using a tri-stage learning rate scheduler with 18K warm-up steps, 36K decay steps, and a peak learning rate of 1e-4. Stage 2 jointly trains ASR, VSR, AVSR, and SLT tasks with the same scheduler structure but 500 warm-up steps and 29.5K decay steps. All experiments are conducted on 8 NVIDIA A6000 GPUs (48GB each), with a maximum of 1,250 video frames processed per GPU per batch. A fixed random seed of 1 is used for reproducibility. We save two checkpoints during training: (i) the best checkpoint based on next-token prediction accuracy on the validation set, and (ii) the last checkpoint at the end of training. The final model is selected by comparing both checkpoints based on BLEU-4 (for SLT) and WER (for ASR, VSR, and AVSR). Due to computational constraints, we report results from a single run. For evaluation, we use beam search (width 5, temperature 0.3) for ASR, VSR, and AVSR tasks, and greedy decoding for SLT.
V Experimental Results
V-A Comparison with the state-of-the-art methods
In Table I, we compare the SLT performance of our unified framework against prior methods on the How2Sign dataset. Compared to PGG-SLT [63], the previous best model without using external data, our approach improves BLEU-4 by +1.5 (15.2 vs. 13.7) and ROUGE by +3.3 (36.2 vs. 32.9). It also outperforms LiTFiC [65] by +1.8 BLEURT (47.1 vs. 45.3). These results highlight the advantage of jointly modeling SLT and VSR within a single framework.
In Table II, we compare performance on VSR, ASR, and AVSR using the LRS3 dataset. Since our model is LLM-based, we focus on comparisons with recent LLM-based approaches, all of which are task-specific. For VSR, our method achieves a WER of 25.9%, comparable to task-specific models such as VSP-LLM and Llama-AVSR. In ASR, our model reaches a WER of 0.79% using only 433 hours of training data, matching the performance of Llama-AVSR, which was trained on 1,759 hours. For AVSR, our model obtains a WER of 0.93%, close to MMS-Llama (0.90%). Compared to USR [14], a recent unified model for VSR, ASR, and AVSR, our method performs better on ASR (0.79% vs. 1.2%) and AVSR (0.93% vs. 1.1%) while using significantly less training data. These results demonstrate that a single unified model can effectively process sign language, lip movements, and audio signals, achieving competitive or superior performance compared to task-specific models across all four tasks. Moreover, the strong performance across tasks suggests that shared linguistic representations and cross-modal learning can generalize effectively across modalities.
TABLE I: Performance comparison of SLT models on the How2Sign. $\dagger$ indicates models that are pretrained on a larger ASL corpus, YouTube-ASL [59], which contains 984 hours of video. “AV Input” indicates whether the model handles audio or lip movement modalities.
@@ V-A Comparison with the state-of-the-art methods
SSLT [73] - - 55.7
Youtube-ASL [59] 12.4 - -
Uni-Sign [74] 14.9 36.0 49.4
SSVP-SLT [6] 15.5 38.4 49.6
SSLT [73] - - 34.0
SSVP-SLT [6] 7.0 25.7 39.3
Fla-LLM [4] 9.7 27.8 -
VAP [5] 12.9 27.8 -
LiTFiC [65] 12.7 32.5 45.3
PGG-SLT [63] 13.7 32.9 -
@@
TABLE II: Performance comparison of VSR, ASR, and AVSR models on LRS3. $\dagger$ indicates models that leverage pseudo-labels generated from the VoxCeleb2 [70] dataset (1,326 hours) during fine-tuning. SL denotes sign language.
@@
V2P [75] 3,886 55.1
RNN-T [12] 31,000 33.6 4.8 4.5
VTP [76] 2,676 30.7
Auto-AVSR [45] 1,902 23.5 1.0 1.0
Auto-AVSR [45] 3,448 19.1 1.0 0.9
ViT3D-CM [77] 90,000 17.0 1.6
SynthVSR [78] 6,720 16.9
LP Conf [79] 100,000 12.8 0.9
@@
AV-HuBERT [43] 1,759 26.9
RAVEn [44] 1,759 23.1 1.4
USR [14] 1,759 22.3 1.2 1.1
AV-HuBERT [43] 433 28.6 1.3 1.4
VATLM [17] 433 28.4 1.2
RAVEn [44] 433 28.2 1.4
BRAVEn [80] 433 26.6 1.2
u-HuBERT [13] 433 29.1 1.5 1.3
@@
Llama-AVSR [20] 1,759 24.0 0.79 0.77
MMS-Llama [18] 1,759 0.72
VSP-LLM [47] 433 25.4
Llama-AVSR [20] 433 25.3 1.1 0.95
MMS-Llama [18] 433 0.90
@@
TABLE III: Ablation study on SLT performance on How2Sign, with progressively added lip encoder and VSR loss.
@@
Figure 2: Multi-task training progress for AVSR, ASR, VSR, and SLT. Curves show next-token prediction accuracy per epoch. Audio-related tasks are plotted in blue; visual-related tasks in red.
TABLE IV: Effect of our two-stage multi-task training strategy. Stage 1 trains only on visual tasks (VSR, SLT), while Stage 2 jointly trains all tasks.
@@
TABLE V: Effect of modality-specific dropout probabilities on multi-task learning performance.
@@
V-B Ablation study
V-B1 Effectiveness of Lip-Aware SLT via VSR Modeling
To assess the effectiveness of our lip-aware SLT framework, which incorporates auxiliary VSR modeling within a unified architecture, we conduct controlled ablation experiments by incrementally adding each component. Four configurations are evaluated, as summarized in Table III.
The base configuration includes the sign encoder, sign length adapter, and LLM decoder, trained solely on the SLT task using the How2Sign dataset. This setup yields a BLEU-4 score of 12.4, ROUGE of 32.6, and BLEURT of 44.6. We then add a lip encoder that processes the signer’s mouth region and injects this representation into the LLM alongside sign features. This addition leads to substantial gains: +1.5 BLEU-4, +2.4 ROUGE, and +1.6 BLEURT (13.9 / 35.0 / 46.2), highlighting the importance of explicitly modeling non-manual cues such as lip movements. Finally, we incorporate a VSR loss to train our framework by leveraging an external LRS3 dataset featuring non-signers. This additional supervision further improves the scores to 14.4 BLEU-4, 35.6 ROUGE, and 46.4 BLEURT. We attribute this improvement to the auxiliary task encourages the LLM to internalize the grammar, vocabulary, and contextual semantics of the spoken target language, thereby enhancing its ability to generate well-formed translations from visual sign inputs. This aligns with findings from recent work [73], which showed that jointly training machine translation with SLT improves overall translation performance.
V-B2 Effectiveness of Two-stage Multi-Task Learning
As shown in Figure 2, preliminary multi-task training experiments revealed significant convergence imbalances: audio-based tasks (ASR, AVSR) quickly surpassed 90% next-token prediction accuracy within 2–3 epochs, while visual tasks (VSR, SLT) improved more slowly. Motivated by these observations, we propose a two-stage multi-task learning strategy. To evaluate the effectiveness of this approach, we conducted three configurations: (1) all tasks trained jointly from scratch; (2) only Stage 1 training; and (3) full two-stage training. Results are summarized in Table IV.
@@ V-B2 Effectiveness of Two-stage Multi-Task Learning
V-B3 Impact of Modality Dropout on Multi-Task Learning
Since the LRS3 dataset provides paired audio-visual inputs, dropping the lip modality corresponds to performing ASR, while dropping the audio modality corresponds to performing VSR. Based on this fact, we leverage modality dropout [43] as an implicit mechanism to control the sampling ratio of each task during training.
@@ V-B3 Impact of Modality Dropout on Multi-Task Learning
TABLE VI: Performance comparison between task-specific training and joint multi-task training using our proposed framework. Each of the first four rows shows the result when the model is trained on a single task only.
@@
V-B4 Task-Specific Training Results
To verify the benefit of multi-task learning in our unified framework, we trained four additional task-specific models: SLT, VSR, ASR, and AVSR, each using only the corresponding task-specific objective. In contrast, the multi-task model is trained using our two-stage multi-task learning strategy. As shown in Table VI, the multi-task model achieves substantial improvements in SLT and ASR, with a BLEU-4 score of 15.2 and a WER of 0.79, respectively. While VSR and AVSR show slightly higher WERs than their task-specific counterparts (by 0.3 and 0.03, respectively), the performance remains comparable, demonstrating that our unified model effectively handles all tasks within a single architecture.
Figure 3: Visualization of attention distributions and qualitative comparisons of SLT predictions with and without lip features. From top to bottom: (1) signer frames, (2) corresponding lip-region crops, (3) attention scores averaged over all heads in the last LLM layer, and (4) ground-truth and predicted sentences. Examples on the left and right correspond to the text tokens “web” and “temperament,” respectively.
V-B5 Qualitative Results
To visually verify the effectiveness of lip-movement modeling in the SLT task, we analyze the attention scores averaged over all heads in the last layer with respect to the LLM input sequence index (x-axis), as shown in Figure 3. We also visualize the input frames corresponding to regions with high attention scores, providing an intuitive comparison between the model’s focus and the visual cues in the sign and lip movements.
On the left side of Figure 3, qualitative examples compare predictions with and without lip features. In the example, the ground-truth sentence is “In the background I have gray pieces of paper with some silver pen that I made the spider web.” Without lip features, the model incorrectly predicts “…that I made the snowflake,” failing to capture the intended meaning. When lip features are included, it correctly outputs “web”. The corresponding high-attention regions align with mouth movements resembling the articulation of “web”. On the right side, another example shows the ground-truth sentence “You need to think about your horse’s temperament.” Without lip features, the model produces the incorrect phrase “…your horse’s depth of,” whereas with lip features it accurately generates “temperament.” These results visually confirm that visual cues from lip movements provide critical articulatory evidence for distinguishing semantically similar expressions.